OpenStudy (anonymous):

square root of 27 a^2b^3c^5

OpenStudy (anonymous):

\[\sqrt{27a ^{2}b ^{3}c ^{5}}\]

OpenStudy (dumbcow):

3abc^2sqrt(3bc)

OpenStudy (anonymous):

27 = 3 times 3 times 3 = 3^2 times 3 b^3 = b^2 times b c^5 = c^2 times c^2 times c

OpenStudy (anonymous):

I know the answer I need an explanation

OpenStudy (anonymous):

so if you do sqrt (27) that is sqrt(3^3 times3) = 3sqrt3 similarly the others too follow the same pattern

OpenStudy (anonymous):

i meant sqrt(3^2 times3)

OpenStudy (anonymous):

ok im clear there what about the variables

OpenStudy (anonymous):

same thing. sqrt (a^2) = a sqrt(b^3) = b sqrt(b) sqrt(c^5) = c^2 sqrt(c)

OpenStudy (anonymous):

what is confusing me is the c^5 i see that c^2c^2 c = c^5

OpenStudy (anonymous):

yes c^2*c^2 = (c^2)^2

OpenStudy (anonymous):

so c^5 = (c^2)^2 times c

OpenStudy (anonymous):

ohhhhhhhhhhhhh your the best thanks so much

OpenStudy (anonymous):

you are welcome.

OpenStudy (anonymous):

for example, 16 = 4^2 but 4 = 2^2 so 4^2 = (2^2)^2 so sqrt 16 = sqrt(4^2) = sqrt (2^2)^2 = 2^2 = 4.